Grade 12Physical Chemistry

The total vapour pressure of a 4mole percent solution of Ammonia in water is 50 torr the vapour pressure of pure water is 17 tort at same temp 293K apply Raults law and Henry law calculate total Vapour pressure of 5mole percent solution

Kirtan shukla
8 Years agoGrade 12

1 Answer

Sandeep Bhamoo
8 Years ago
Let vapour pressure of pure ammonia=x, then ,50=17(0.96) +x(0.04), x=842torr. Now new mole fraction are 0.95(water), 0.05(ammonia). Hence P= 0.95(17) +0.05(842) =58.25 torr. Please do approve my answer if correct